A 0.75-μF capacitor is connected to a 400-V battery. Find the charge on the capacitor.

5.3E+4 C

3.0E–4 C

5.3E–8 C

1.9E–3 C

4.0E–2 C
Answer #1

Given, capacitance, C = 0.75 F = 0.75*10-6

voltage, V = 400 volt

Thus, charge stored, Q = CV = (0.75*10-6*400) coulomb

= 3*10-4 Coulomb

So, option (b) 3.0E-4 C.